# Lifting subgroups of $\mathrm{PSL}_{2}$ to $\mathrm{SL}_{2}$ over local fields

### v1

@@ 5. Proof of Propositions D and E
Given a representation $\rho:\Gamma\to{\rm PSL_{2}}(K)$, we can pull it back along the quotient $\pi:{\rm SL_{2}}(K)\to{\rm PSL_{2}}(K)$ to get a representation $\rho^{\prime}:\Gamma^{\prime}\to{\rm SL_{2}}(K)$, with $\Gamma^{\prime}$ a central extension of $\Gamma$ by $\mathbb{Z}/2\mathbb{Z}$. The situation is as in this commutative digram; both vertical rows are the middle three terms of a central extension.
{tikzcd}
@@
We follow the same argument as [3, Proposition 2]. First note that $A_{n}$ is simple for $n\geq 5$ and $H^{2}(A_{n};\mathbb{Z}/2\mathbb{Z})=\mathbb{Z}/2\mathbb{Z}\neq 0$ for $n\geq 4$, which follows from their Schur multipliers [13, page 27] and a short calculation with the Universal Coefficient Theorem [6, Theorem 3.2]. Furthermore, for each $n\geq 6$, the group $A_{n}$ does not occur as a subgroup of ${\rm PSL_{2}}(K)$, except for $A_{6}\cong{\rm PSL_{2}(9)}$ when $\mathrm{char}(K)=3$. Indeed, the only feasible possibilities for an isomorphism between $A_{n}$ and the subgroups of ${\rm PSL_{2}}(K)$ listed in [11, Theorem 1] are with a group of Borel type, or a group of the form ${\rm PGL}(2,q^{\prime})$ or ${\rm PSL}(2,q^{\prime})$ for some divisor $q^{\prime}$ of $q$, all of which occur only in positive characteristic. The first two possibilities would contradict that $A_{n}$ is simple and, other than the case $A_{6}\cong{\rm PSL_{2}}(9)$, the last possibility is ruled out by [13, page 3]. Hence, when $n\geq 7$ or $n=6$ and $\mathrm{char}(K)\neq 3$, the only representation of $A_{n}$ into ${\rm PSL_{2}}(K)$ is the identity representation. This can certainly be lifted, which completes the proof. ∎

### v2

@@ 5. Proof of Propositions D and E
Given a representation $\rho:\Gamma\to{\rm PSL_{2}}(K)$, we can pull it back along the quotient $\pi:{\rm SL_{2}}(K)\to{\rm PSL_{2}}(K)$ to get a representation $\rho^{\prime}:\Gamma^{\prime}\to{\rm SL_{2}}(K)$, with $\Gamma^{\prime}$ a central extension of $\Gamma$ by $\mathbb{Z}/2\mathbb{Z}$. The situation is as in this commutative diagram; both vertical rows are the middle three terms of a central extension.
[…]
@@
We follow the same argument as [3, Proposition 2]. First note that $A_{n}$ is simple for $n\geq 5$ and $H^{2}(A_{n};\mathbb{Z}/2\mathbb{Z})=\mathbb{Z}/2\mathbb{Z}\neq 0$ for $n\geq 4$, which follows from their Schur multipliers [15, page 27] and a short calculation with the Universal Coefficient Theorem [7, Theorem 3.2]. Furthermore, for each $n\geq 6$, the group $A_{n}$ does not occur as a subgroup of ${\rm PSL_{2}}(K)$, except for $A_{6}\cong{\rm PSL_{2}(9)}$ when $\mathrm{char}(K)=3$. Indeed, the only feasible possibilities for an isomorphism between $A_{n}$ and the subgroups of ${\rm PSL_{2}}(K)$ listed in [13, Theorem 1] are with a group of Borel type, or a group of the form ${\rm PGL}(2,q^{\prime})$ or ${\rm PSL}(2,q^{\prime})$ for some divisor $q^{\prime}$ of $q$, all of which occur only in positive characteristic. The first two possibilities would contradict that $A_{n}$ is simple and, other than the case $A_{6}\cong{\rm PSL_{2}}(9)$, the last possibility is ruled out by [15, page 3]. Hence, when $n\geq 7$ or $n=6$ and $\mathrm{char}(K)\neq 3$, the only representation of $A_{n}$ into ${\rm PSL_{2}}(K)$ is the identity representation. This can certainly be lifted, which completes the proof. ∎